Here is an example of a problem that I am having a hard time figuring out. Given f(x) = 2-2x and g(x) = -6x+2, find f+g
Answer by Alan3354(69443)

Here is an example of a problem that I am having a hard time figuring out. Given f(x) = 2-2x and g(x) = -6x+2, find f+g
------------
f(?) + g(?) ??? you have to specify.
If it's of x, then just add them.
f(x) + g(x) = 2-2x + (-6x+2)
= -8x + 4